# Interacting binaries on the Main Sequence as in-situ tracers of mass transfer efficiency and stability

### v1

@@ Abstract
Understanding the transfer of mass and angular momentum in binary interactions is crucial for modelling the evolution of any interacting binary after the first mass transfer phase. Our assumptions on mass transfer physics shape the predictions for later stages of binary evolution, such as the immediate progenitors of stripped-envelope supernovae and gravitational wave mergers. We constrain the efficiency and stability of thermal timescale mass transfer in massive binary evolution using the observed population of massive interacting binaries on the Main Sequence (‘Algols’) in the Milky Way, Large and Small Magellanic Clouds. Assuming the present-day mass of the donor star represents its initial convective core mass at Zero-Age Main Sequence, we estimate its initial mass using detailed stellar evolution models. From the initial donor mass, we calculate the range of initial accretor masses (for different mass transfer efficiencies). By imposing physical constraints on the above initial parameter ranges, we derive the mass transfer efficiency, stability and angular momentum loss that can reproduce the current properties of each Algol binary. We find that purely conservative or non-conservative mass transfer cannot explain the current mass ratio and orbital period of all massive Algols. Angular momentum conservation rules out conservative mass transfer in $\sim$28 % of massive Algols in the SMC. About three-quarters of all massive Algols are consistent with having undergone inefficient mass transfer ($\lesssim$ 50 %), while the remaining systems, mostly residing in the LMC and Milky Way, require mass transfer to have been more efficient than 25%. For our fiducial assumption on the extent of envelope stripping, the current sample of massive Algols does not require mass transfer to be efficient at the shortest orbital periods (2 d) at any metallicity. We find evidence that mass transfer on the Main Sequence needs to be stable for initial accretor-to-donor mass ratios as unequal as $\sim 0.6$. Unless biased by observational selection effects, the massive Algols in the SMC seem to have undergone less efficient mass transfer than those in the LMC and Milky Way.
@@ 1 Introduction
Massive stars are rarely found alone (D. Vanbeveren et al., 1998b; H. A. Kobulnicky & C. L. Fryer, 2007; B. D. Mason et al., 2009; H. Sana et al., 2012; M. Moe & R. Di Stefano, 2017; S. S. R. Offner et al., 2023). Surveys of O- and B-type stars have established that a majority reside in multiples (H. A. Abt et al., 1990; D. C. Kiminki et al., 2008; L. Mahy et al., 2009, 2013; R. Chini et al., 2012; H. A. Kobulnicky et al., 2012, 2014; A. Sota et al., 2014; P. R. Dunstall et al., 2015; H. Sana et al., 2008, 2009, 2011, 2013, 2014; R. H. Barbá et al., 2017; L. A. Almeida et al., 2017; E. Trigueros Páez et al., 2021; J. I. Villaseñor et al., 2021, 2025; T. Shenar et al., 2022, 2024; G. Banyard et al., 2022; B. W. Ritchie et al., 2022; M. Ramírez-Tannus et al., 2024). Most of them are so close that interactions through tides (J. P. Zahn, 1977; P. Goldreich & P. D. Nicholson, 1989; P. Marchant et al., 2016; I. Mandel & S. E. de Mink, 2016; S. E. de Mink & I. Mandel, 2016; M. Sun et al., 2023; S. Rosu, 2024; K. A. Rocha et al., 2025; K. Sen et al., 2025), colliding winds (G. Rauw & Y. Nazé, 2016; A. Kashi, 2020; A. Kashi et al., 2022; Y. Nazé et al., 2022), Roche-lobe overflow (B. Paczyński, 1971; P. Podsiadlowski et al., 1992; D. Vanbeveren et al., 1998a; N. Langer et al., 2003), and/or mergers (P. Podsiadlowski et al., 2006; F. R. N. Schneider et al., 2019; S. Wu et al., 2020; A. Menon et al., 2024; R. A. Patton et al., 2025) are inevitable. These interactions profoundly shape the evolutionary pathways of massive stars (S. E. de Mink et al., 2014; K. Sen et al., 2023), their feedback on the surrounding medium (J. J. Eldridge et al., 2008; E. R. Stanway et al., 2016; Y. Götberg et al., 2017), and their ultimate fates as progenitors of stripped-envelope supernovae (SESNe, S. C. Yoon et al., 2010; E. Zapartas et al., 2017; N. Sravan et al., 2020; E. Laplace et al., 2020, 2021; D. R. Aguilera-Dena et al., 2022, 2023; A. Ercolino et al., 2024, 2025) and compact object binaries that may later merge as gravitational-wave (GW) sources (K. Belczynski et al., 2008; J. J. Eldridge & E. R. Stanway, 2016; M. U. Kruckow et al., 2018; M. Spera et al., 2019; F. S. Broekgaarden et al., 2022; L. A. C. van Son et al., 2022; J. Riley et al., 2022; M. M. Briel et al., 2023). Understanding the physical processes at play in these systems is therefore essential not only for stellar astrophysics (P. Marchant & J. Bodensteiner, 2024) but also for cosmology (P. F. Hopkins et al., 2014) and predicting the rates and properties of transient populations across the Universe (N. Smith et al., 2011; I. Mandel & F. S. Broekgaarden, 2022).
@@
However, the treatment of mass transfer and angular momentum loss in massive binaries remains one of the most significant sources of uncertainty in stellar evolution theory (M. S. Hjellming & R. F. Webbink, 1987; G. E. Soberman et al., 1997; T. E. Woods & N. Ivanova, 2011; K. Pavlovskii et al., 2017; H. Ge et al., 2010, 2015, 2020; K. Temmink et al., 2023; C. Schürmann & N. Langer, 2024). Whether the mass transfer is highly efficient or largely inefficient (C. A. Nelson & P. P. Eggleton, 2001; J. Petrovic et al., 2005; S. E. de Mink et al., 2007; Y. Shao & X.-D. Li, 2016; S. Vinciguerra et al., 2020; Y. Bouffanais et al., 2021; K. Sen et al., 2022; I. Romero-Shaw et al., 2023; M. Nuijten & G. Nelemans, 2025; T. Lechien et al., 2025; C. Schürmann et al., 2025), and how the lost mass carries away angular momentum (M. MacLeod et al., 2018a, b; W. Lu et al., 2023), dramatically alters the orbital period and mass-ratio evolution of binaries (J. J. Eldridge et al., 2011; M. Renzo et al., 2019; R. Willcox et al., 2023; T. Wagg et al., 2025). Population synthesis studies demonstrate that these choices can alter the diversity and rates of SESNe and GW mergers respectively (J. J. Eldridge et al., 2013; P. Agrawal et al., 2022, 2023; K. Belczynski et al., 2022; A. P. Boesky et al., 2024; L. Rauf et al., 2024; E. Zapartas et al., 2025; D. Souropanis et al., 2025), while also reshaping the predicted orbital period distribution of compact object binaries, and masses and spin magnitudes of compact objects (K. Belczynski et al., 2020; A. Olejak et al., 2021, 2024; T. M. Tauris, 2022; M. Y. M. Lau et al., 2024; A. Dorozsmai & S. Toonen, 2024; H. C. G. Larsen et al., 2025). Quantitatively constraining these processes is thus critical for linking present-day massive binaries with their end states (e.g. I. Mandel, 2025). Without empirical constraints on early stages of binary evolution, parameter uncertainties and degeneracies in the predicted population of SESNe or GW mergers may not be resolved simply by an increase in their observed number in the coming decades.
@@
An empirical approach is to study Algol-type binaries (L. P. Surkova & M. A. Svechnikov, 2004; S. E. de Mink et al., 2007; N. Mennekens & D. Vanbeveren, 2017; O. Y. Malkov, 2020; K. Sen et al., 2022), which are interacting systems where mass transfer occurs while both stars are still on the Main Sequence (‘Case A’, R. Kippenhahn & A. Weigert, 1967; O. R. Pols, 1994; C. A. Nelson & P. P. Eggleton, 2001). These systems are especially valuable because they represent the earliest phase of mass transfer, prior to the additional complexities introduced by envelope expansion (A. Romagnolo et al., 2023), strong winds (N. Smith, 2014; M. Renzo et al., 2017; J. Josiek et al., 2024), or advanced nuclear burning (A. Heger et al., 2002; R. Farmer et al., 2016; A. Grichener et al., 2025). As such, they require the fewest assumptions regarding prior binary evolution, offering a relatively clean probe of the efficiency and stability of thermal-timescale mass transfer.
@@
In Case A binaries, mass transfer begins on the donor’s thermal timescale (fast Case A) and transitions into a slower, nuclear-timescale phase as the donor expands during core hydrogen burning. This long-lived ‘slow Case A’ stage, which spans millions of years, yields observable semi-detached systems. Since all such binaries must have previously undergone an intense, fast Case A episode, comparing models with observed Algols constrains both phases. Moreover, Case A is especially relevant at high masses, where it is more common (S. E. de Mink et al., 2007; K. Sen et al., 2023; C. A. Burt et al., 2025), making massive Algols bright, numerous, and ideal testbeds for mass transfer physics. The demographics of observed Algols across different metallicities-Milky Way, Large Magellanic Cloud (LMC) and Small Magellanic Cloud (SMC)-therefore provide a powerful benchmark against which theoretical models can be tested (S. E. de Mink et al., 2007; K. Sen et al., 2022).
@@
In this work, we study the mass transfer efficiency, angular momentum loss, and stability of mass transfer in the observed populations of massive Algols in the Milky Way, LMC and SMC, and investigate any metallicity difference. We develop an analytical framework that links current observable binary properties—orbital period, donor mass, and accretor mass—to their initial parameters. By imposing physically motivated boundaries on the initial parameters, we derive constraints on the range of mass transfer efficiency and angular momentum loss during the mass transfer episode that can reproduce the observed population. We also determine limits on the stability of mass transfer required to obtain meaningful solutions for mass transfer efficiency.
@@
This paper is organised as follows. Section 2 describes the observational sample and our methodology for deriving initial binary parameters. Section 3 presents the range of mass transfer efficiency and angular momentum loss possible, for individual systems to overall trends across the Algol population. In Sect. 4, we compare our results with the literature and evaluate the robustness of our assumptions. We summarise our main conclusions in Sect. 5.
2 Methods
2.1 Observed systems
@@
In this study, we include all the massive Algol binaries observed in the Magellanic Clouds and the Milky Way. In the SMC, we use the catalogue in S. E. de Mink et al. (2007, Table 1, 29 systems) curated from the observational studies of the T. J. Harries et al. (2003) and R. W. Hilditch et al. (2005). We take the massive Algols listed in Table 1 and Table 2 of K. Sen et al. (2022, 33 systems) in the LMC and the Milky Way, respectively (heterogeneous sample, see references within for details on individual systems). We list the current masses of the donors $M_{\rm d}$, accretors $M_{\rm a}$ and binary orbital periods $P_{\rm orb}$ used in this study for all the systems in Table 1 and Table 2.
@@
2.2 Initial stellar and binary parameters
@@
For each value of initial donor mass, there exists a maximum initial orbital period $P_{\rm orb,i,max}$ up to which mass transfer can initiate on the Main Sequence (C. A. Nelson & P. P. Eggleton, 2001; S. E. de Mink et al., 2007; K. Sen et al., 2022; D. Pauli et al., 2022). This maximum initial orbital period is weakly dependent on the initial mass ratio of the binary, but sensitive to stellar radii and the physics that determines them (opacity primarily, and at second order mass loss and inflation, e.g., D. Sanyal et al. 2015; C. Xin et al. 2022). For each massive Algol, we estimate the maximum initial orbital period $P_{\rm orb,i,max}$ that the binary could have started with, using the binary models described in Sect. 2.4 (listed in Table 1) and Table 2.
The difference between the initial donor mass and the current donor mass in each Algol donor gives the mass lost by the donor $\Delta M_{\rm d}$ and is equal to the maximum amount of mass that the mass accreting star could have accreted during the fast thermal timescale mass transfer phase. Substracting the above difference from the present mass of the accretor gives the lowest initial mass of the accretor $M_{\rm a,i,min}$ at its Zero-Age Main Sequence, if the accretor could accrete all the transferred mass (conservative mass-transfer limit). Here, we limit the lowest possible initial accretor mass to be 0.08 $M_{\odot}$, required to initiate core hydrogen burning in its core.
@@
2.3 Mass transfer efficiency and specific angular momentum loss
@@
Equation (1) can be written in terms of the total mass $M_{\rm T}=M_{\rm a}+M_{\rm d}$ and mass ratio $q=M_{\rm a}/M_{\rm d}$ of the binary system as (G. E. Soberman et al., 1997; M. Nuijten & G. Nelemans, 2025)
@@
For conservative mass transfer $\varepsilon=1$, there is no angular momentum lost from the binary (neglecting stellar winds; moreover, winds tap into the spin angular momentum of the stars, which is much smaller than the orbital angular momentum), and the initial orbital period $P_{\rm ini,cons}$ is given by (e.g. T. M. Tauris & E. P. J. van den Heuvel, 2023)
@@
where $\gamma\,(>0)$ is a multiplicative factor to capture different modes of mass loss (O. R. Pols & M. Marinus, 1994), $J_{\rm orb}$ is the total orbital angular momentum of the binary, $j_{\rm orb}$ is the specific orbital angular momentum of the binary and $j_{\rm loss}$ is the specific angular momentum lost from the binary.
In reality, $\gamma$ depend on how the mass flows away from the binary (e.g. G. E. Soberman et al., 1997; W. Lu et al., 2023), and thus could vary as the stars evolve through mass transfer. As a first approximation, we consider $\gamma$ as constant to relate analytically the initial and present-day period. The initial orbital period $P_{\rm orb,i}$ can be related to the orbital period and masses of the binary components (see Eq. (A2) of O. R. Pols & M. Marinus, 1994)
@@
We calculate the initial orbital period values for $\gamma$ = 0, 1, 2, 3. Values of $\gamma>3$ are expected to be unlikely (see discussion in Sect. 2.4 of M. Nuijten & G. Nelemans, 2025). Magnetic braking (D. A. Bour et al., 2025) or L2 mass loss (W. Lu et al., 2023) can lead to high values of $\gamma$, but most massive stars are not found to host detectable dipolar magnetic fields (G. A. Wade et al., 2016). The strong differential rotation during the accretion phase, however, may generate magnetic fields (J. Braithwaite & H. C. Spruit, 2017). For isotropic re-emission from the surface of the accretor, $\gamma=M_{\rm d}/M_{\rm a}$. We discuss the possibility of $\gamma>3$ in Sect. 4. The initial orbital period $P_{\rm orb,i}$ is limited by the maximum initial orbital period $P_{\rm orb,i,max}$ up to which mass transfer can initiate on the Main Sequence. The initial orbital period (for $\varepsilon$ constant over time, Eq. (A4-A5) of O. R. Pols & M. Marinus, 1994) is given by
$\ln\left(\frac{P_{\rm orb,i}}{P_{\rm orb}}\right)=3\ln\left(\frac{M_{\rm d}}{M_{\rm d,i}}\right)+\frac{3}{\varepsilon}\ln\left(\frac{M_{\rm a}}{M_{\rm a,i}}\right)+\frac{3}{2}\ln\left(\frac{M_{\rm T}}{M_{\rm T,i}}\right),\rm{for\,\,}0<\varepsilon\leq 1$ (6)
@@
2.4 Detailed binary evolution models

### v2

@@ Abstract
Understanding the transfer of mass and angular momentum in binary interactions is crucial for modelling the evolution of any interacting binary after the first mass transfer phase. Mass transfer physics assumptions shape the predictions for later stages of binary evolution, such as the immediate progenitors of stripped-envelope supernovae and gravitational wave mergers. We constrain the efficiency and stability of thermal timescale mass transfer in massive binary evolution using the observed population of 62 massive interacting binaries on the Main Sequence (‘Algols’) in the Milky Way, Large and Small Magellanic Clouds. We find that purely conservative or non-conservative mass transfer cannot explain the current mass ratio and orbital period of all massive Algols. Angular momentum conservation rules out conservative mass transfer in $\sim$28 % of massive Algols in the SMC. About three-quarters of all massive Algols are consistent with having undergone inefficient mass transfer ($\lesssim$ 50 %), while the remaining systems, mostly residing in the LMC and Milky Way, require mass transfer to have been more efficient than 25%. For our fiducial assumption on the extent of envelope stripping, the current sample of massive Algols does not require mass transfer to be efficient at the shortest orbital periods ($\sim$2 d) at any metallicity. We find evidence that mass transfer on the Main Sequence needs to be stable for initial accretor-to-donor mass ratios as unequal as $\sim 0.6$. Unless biased by observational selection effects, the massive Algols in the SMC seem to have undergone less efficient mass transfer than those in the LMC and Milky Way.
@@ I Introduction
Massive stars ($\geq 8\,M_{\odot}$) are rarely found alone (Vanbeveren et al., 1998b; Kobulnicky and Fryer, 2007; Mason et al., 2009; Sana et al., 2012, 2014; Moe and Di Stefano, 2017; Offner et al., 2023). Surveys of O- and B-type stars have established that a majority reside in multiples (Abt et al., 1990; Kiminki et al., 2008; Mahy et al., 2009, 2013; Kobulnicky et al., 2012, 2014; Sota et al., 2014; Dunstall et al., 2015; Sana et al., 2008, 2009, 2011, 2013, 2014; Barbá et al., 2017; Almeida et al., 2017; Trigueros Páez et al., 2021; Villaseñor et al., 2021, 2025; Shenar et al., 2022, 2024; Banyard et al., 2022; Ritchie et al., 2022; Ramírez-Tannus et al., 2024; Frost et al., 2025). Most of them are so close that interactions through tides (Zahn, 1977; Goldreich and Nicholson, 1989; Marchant et al., 2016; Mandel and de Mink, 2016; de Mink and Mandel, 2016; Sun et al., 2023; Rosu, 2024; Rocha et al., 2025; Sen et al., 2025), colliding winds (Rauw and Nazé, 2016; Kashi, 2020; Kashi et al., 2022; Nazé et al., 2022), Roche-lobe overflow (Paczyński, 1971; Podsiadlowski et al., 1992; Vanbeveren et al., 1998a; Langer et al., 2003), and/or mergers (Podsiadlowski et al., 2006; Schneider et al., 2019; Wu et al., 2020; Menon et al., 2024; Frost et al., 2024; Patton et al., 2025) are inevitable. These interactions profoundly shape the evolutionary pathways of massive stars (de Mink et al., 2014; Sen et al., 2023; Jin et al., 2025), their feedback on the surrounding medium (Eldridge et al., 2008; Stanway et al., 2016; Götberg et al., 2017), and their ultimate fates as progenitors of stripped-envelope supernovae (SESNe, Yoon et al., 2010; Zapartas et al., 2017; Sravan et al., 2020; Laplace et al., 2020, 2021; Aguilera-Dena et al., 2022, 2023; Ercolino et al., 2024, 2025) and compact object binaries that may later merge as gravitational-wave (GW) sources (Belczynski et al., 2008; Eldridge and Stanway, 2016; Kruckow et al., 2018; Spera et al., 2019; Broekgaarden et al., 2022; van Son et al., 2022; Riley et al., 2022; Briel et al., 2023). Understanding the physical processes at play in these systems is therefore essential not only for stellar astrophysics (Marchant and Bodensteiner, 2024) but also for cosmology (Hopkins et al., 2014) and predicting the rates and properties of transient populations across the Universe (Smith et al., 2011; Mandel and Broekgaarden, 2022).
@@
However, the treatment of mass transfer and angular momentum loss in massive binaries remains one of the most significant sources of uncertainty in stellar evolution theory (Hjellming and Webbink, 1987; Soberman et al., 1997; Woods and Ivanova, 2011; Pavlovskii et al., 2017; Ge et al., 2010, 2015, 2020; Temmink et al., 2023; Schürmann and Langer, 2024). Whether the mass transfer is highly efficient or largely inefficient (Nelson and Eggleton, 2001; Petrovic et al., 2005; de Mink et al., 2007; Shao and Li, 2016; Vinciguerra et al., 2020; Bouffanais et al., 2021; Sen et al., 2022; Romero-Shaw et al., 2023; Nuijten and Nelemans, 2025; Lechien et al., 2025; Schürmann et al., 2025), and how the lost mass carries away angular momentum (MacLeod et al., 2018a, b; Lu et al., 2023), dramatically alters the orbital period and mass-ratio evolution of binaries (Eldridge et al., 2011; Renzo et al., 2019; Willcox et al., 2023; Wagg et al., 2025). Population synthesis studies demonstrate that these choices can alter the diversity and rates of SESNe and GW mergers respectively (Eldridge et al., 2013; Agrawal et al., 2022, 2023; Belczynski et al., 2022; Boesky et al., 2024; Rauf et al., 2024; Zapartas et al., 2025; Souropanis et al., 2025; Willcox et al., 2025), while also reshaping the predicted orbital period distribution of compact object binaries, and masses and spin magnitudes of compact objects (Belczynski et al., 2020; Olejak et al., 2021, 2024; Tauris, 2022; Lau et al., 2024; Dorozsmai and Toonen, 2024; Larsen et al., 2025). Quantitatively constraining these processes is thus critical for linking present-day massive binaries with their end states (e.g. Mandel, 2025). Without empirical constraints on early stages of binary evolution, parameter uncertainties and degeneracies in the predicted population of SESNe or GW mergers may not be resolved simply by an increase in their observed number in the coming decades.
@@
An empirical approach is to study Algol-type binaries (Surkova and Svechnikov, 2004; de Mink et al., 2007; Mennekens and Vanbeveren, 2017; Malkov, 2020; Sen et al., 2022), which are interacting systems where mass transfer occurs while both stars are still on the Main Sequence (‘Case A’, Kippenhahn and Weigert, 1967; Pols, 1994; Nelson and Eggleton, 2001). These systems are especially valuable because they represent the earliest phase of mass transfer, prior to the additional complexities introduced by envelope expansion (Romagnolo et al., 2023), strong winds (Smith, 2014; Renzo et al., 2017; Josiek et al., 2024), or advanced nuclear burning (Heger et al., 2002; Farmer et al., 2016; Grichener et al., 2025). As such, they require the fewest assumptions regarding prior binary evolution, offering a relatively clean probe of the efficiency and stability of thermal-timescale mass transfer.
@@
In Case A binaries, mass transfer begins on the donor’s thermal timescale (fast Case A) and transitions into a slower, nuclear-timescale phase as the donor expands during core hydrogen burning. This long-lived ‘slow Case A’ stage, which spans millions of years, yields observable semi-detached systems. Since all such binaries must have previously undergone an intense, fast Case A episode, comparing models with observed Algols constrains both phases. Moreover, Case A is especially relevant at high masses, where it is more common (de Mink et al., 2007; Sen et al., 2023; Burt et al., 2025), making massive Algols bright, numerous, and ideal testbeds for mass transfer physics. The demographics of observed Algols across different metallicities-Milky Way, Large Magellanic Cloud (LMC) and Small Magellanic Cloud (SMC)-therefore provide a powerful benchmark against which theoretical models can be tested (de Mink et al., 2007; Sen et al., 2022).
@@
In this work, we study the mass transfer efficiency, angular momentum loss, and stability of mass transfer in the observed populations of 62 massive Algols in the Milky Way, LMC and SMC, and investigate any metallicity difference. We develop an analytical framework that links current observable binary properties—orbital period, donor mass, and accretor mass—to their initial parameters. By imposing physically motivated boundaries on the initial parameters, we derive constraints on the range of mass transfer efficiency and angular momentum loss during the mass transfer episode that can reproduce the observed population. We also determine limits on the stability of mass transfer required to obtain meaningful solutions for mass transfer efficiency.
@@
This paper is organised as follows. Section II describes the observational sample and our methodology for deriving initial binary parameters. Section III presents the range of mass transfer efficiency and angular momentum loss possible, for individual systems to overall trends across the Algol population. In Sect. IV, we compare our results with the literature and evaluate the robustness of our assumptions. We summarise our main conclusions in Sect. V.
II Methods
II.1 Observed systems
@@
In this study, we include all the massive Algol binaries observed in the Magellanic Clouds and the Milky Way. In the SMC, we use the catalogue in de Mink et al. (2007, Table 1, 29 systems) curated from the observational studies of the Harries et al. (2003) and Hilditch et al. (2005). We take the massive Algols listed in Table 1 and Table 2 of Sen et al. (2022, 33 systems) in the LMC and the Milky Way, respectively (heterogeneous sample, see references within for details on individual systems). We list the current masses of the donors $M_{\rm d}$, accretors $M_{\rm a}$ and binary orbital periods $P_{\rm orb}$ used in this study for all the systems in Table 2 and Table 3.
@@
II.2 Initial stellar and binary parameters
@@
For each value of initial donor mass, there exists a maximum initial orbital period $P_{\rm orb,i,max}$ up to which mass transfer can initiate on the Main Sequence (Nelson and Eggleton, 2001; de Mink et al., 2007; Sen et al., 2022; Pauli et al., 2022). This maximum initial orbital period is weakly dependent on the initial mass ratio of the binary, but sensitive to stellar radii and the physics that determines them (opacity primarily, and at second order mass loss and inflation, e.g., Sanyal et al. 2015; Xin et al. 2022). For each massive Algol, we estimate the maximum initial orbital period $P_{\rm orb,i,max}$ that the binary could have started with, using the binary models described in Sect. II.4 (listed in Table 2) and Table 3.
The difference between the initial donor mass and the current donor mass in each Algol donor gives the mass lost by the donor $\Delta M_{\rm d}$ and is equal to the maximum amount of mass that the mass accreting star could have accreted during the fast thermal timescale mass transfer phase. Subtracting the above difference from the present mass of the accretor gives the lowest initial mass of the accretor $M_{\rm a,i,min}$ at its Zero-Age Main Sequence, if the accretor could accrete all the transferred mass (conservative mass-transfer limit). Here, we limit the lowest possible initial accretor mass to be 0.08 $M_{\odot}$, required to initiate core hydrogen burning in its core.
@@
II.3 Mass transfer efficiency and specific angular momentum loss
@@
Equation (1) can be written in terms of the total mass $M_{\rm T}=M_{\rm a}+M_{\rm d}$ and mass ratio $q=M_{\rm a}/M_{\rm d}$ of the binary system as (Soberman et al., 1997; Nuijten and Nelemans, 2025)
@@
For conservative mass transfer $\varepsilon=1$, there is no angular momentum lost from the binary (neglecting stellar winds; moreover, winds tap into the spin angular momentum of the stars, which is much smaller than the orbital angular momentum), and the initial orbital period $P_{\rm orb,i,cons}$ is given by (e.g. Tauris and van den Heuvel, 2023)
@@
where $\gamma\,(>0)$ is a multiplicative factor to capture different modes of mass loss (Pols and Marinus, 1994), $J_{\rm orb}$ is the total orbital angular momentum of the binary, $j_{\rm orb}$ is the specific orbital angular momentum of the binary and $j_{\rm loss}$ is the specific angular momentum lost from the binary.
In reality, $\gamma$ depend on how the mass flows away from the binary (e.g. Soberman et al., 1997; Lu et al., 2023), and thus could vary as the stars evolve through mass transfer. As a first approximation, we consider $\gamma$ as constant to relate analytically the initial and present-day period. The initial orbital period $P_{\rm orb,i}$ can be related to the orbital period and masses of the binary components (see Eq. (A2) of Pols and Marinus, 1994)
@@
We calculate the initial orbital period values for $\gamma$ = 0, 1, 2, 3. Values of $\gamma>3$ are expected to be unlikely (see discussion in Sect. 2.4 of Nuijten and Nelemans, 2025). Magnetic braking (Bour et al., 2025) or L2 mass loss (Lu et al., 2023) can lead to high values of $\gamma$, but most massive stars are not found to host detectable dipolar magnetic fields (Wade et al., 2016). The strong differential rotation during the accretion phase, however, may generate magnetic fields (Braithwaite and Spruit, 2017) and have been recently detected in a few semi-detached and contact binaries (Hubrig et al., 2026). For isotropic re-emission from the surface of the accretor, $\gamma=M_{\rm d}/M_{\rm a}$. We discuss the possibility of $\gamma>3$ in Sect. IV. The initial orbital period $P_{\rm orb,i}$ is limited by the maximum initial orbital period $P_{\rm orb,i,max}$ up to which mass transfer can initiate on the Main Sequence. The initial orbital period (for $\varepsilon$ constant over time, Eq. (A4-A5) of Pols and Marinus, 1994) is given by
$\ln\left(\frac{P_{\rm orb,i}}{P_{\rm orb}}\right)=3\ln\left(\frac{M_{\rm d}}{M_{\rm d,i}}\right)+\frac{3}{\varepsilon}\ln\left(\frac{M_{\rm a}}{M_{\rm a,i}}\right)+\frac{3}{2}\ln\left(\frac{M_{\rm T}}{M_{\rm T,i}}\right),0<\varepsilon\leq 1$ (6)
@@
Table 1: Summary of grid parameters. Notes. $\Delta$ = grid spacing; $N_{\rm Case\,A}$ = number of models undergoing Case A mass transfer. Models here.
Galaxy metallicity initial donor mass $M_{\rm d,i}$ initial orbital period $P_{\rm i}$ initial mass ratio $q_{\rm i}$ $j_{\rm loss}$ $N_{\rm Case\,A}$
[…]
II.4 Detailed binary evolution models